# Life-Code: Central Dogma Modeling with Multi-Omics Sequence Unification

### v2

@@ 3.2 Comparison on Genomic Benchmarks
As DNA sequences are the foundation of the central dogma, we primarily compare with the latest DNA foundation models with the SFT evaluation protocol proposed by (nips2024hyenadna, ; iclr2024dnabert2, ), following benchmark evaluations for fair comparison. We consider three types of DNA model architectures: (a) classical BERT encoders, including DNABERT (BioInfo2021dnabert, ), DNABERT2 (iclr2024dnabert2, ), GENA-LM (fishman2023GenaLM, ), Nucleotide Transformer (NT) (NM2023NucleotideTrans, ), with different tokenizers like VQDNA (icml2024vqdna, ) and MxDNA (nips2024MXDNA, ); (b) RNN-based linear attentions including HyenaDNA (nips2024hyenadna, ) and Caduceus (icml2024caduceus, ); (c) CNN model ConvNova (iclr2025convnova, ).
@@
We first compare Life-Code with 8 evaluation tasks in the Genomic Benchmark (BMC2023genomicbenchmark, ), which we adopted results from GenBench implementations (liu2024genbench, ), and all compared DNA models are fully SFT and reported the top-1 accuracy. Table 1 demonstrates that Life-Code surpasses previous models with remarkable margins since its hybrid designs combine the advantages.
GUE Benchmarks.
We then evaluate Life-Code with 24 long-range genomic tasks on the GUE benchmark (iclr2024dnabert2, ), as shown in Table 2, where 7 widely adopted practical genomic tasks are conducted, including Epigenetic Mark Prediction (EMP) for Yeast, Transcription Factor Prediction on Mouse and Human genome, Promoter Detection, Core Promoter Detection, Splice Site Prediction, and Covid Variants Classification. All results are adopted from DNABERT2 or each paper with their official SFT setups, and the Matthews Correlation Coefficient (MCC) is reported as the metric.
3.3 Comparison on RNA Splicing Site Prediction
Since precursor messenger RNAs (pre-mRNAs) splicing is a crucial process in eukaryotic gene expression, we then evaluate Life-Code on Splicing Site Prediction tasks as binary classification tasks. We consider two typical datasets with the SFT evaluation protocol. As for the SpliceAI dataset (JAGANATHAN2019SpliceAI, ), we evaluate DNA models with top-1 AUC scores and find that Life-Code consistently outperforms popular SpliceAI and DNA models, indicating its great ability to model the central dogma. As for Spliceator dataset (BMC2021spliceator, ), we then consider RNA splicing methods (Spliceator, DDSP (naito2019DDSP, ), RINALMo (Penic2024RiNALMo, )) and RNA language model, RNA-FM (shen2024rnafm, ), with RNA data of multi-species. As shown in Table 4, our Life-Code achieves the best F1 scores with the specified baselines.
@@ 3.3 Comparison on RNA Splicing Site Prediction
Table 4: Comparison on mRNA Splicing Site Prediction with two splicing datasets from SpliceAI and Spliceator. The top-1 AUC score is reported.
@@ 3.4 Comparison on Protein Fitness Prediction
Following EVO (nguyen2024evo, ), we evaluate foundation models in a strict zero-shot setting on Protein Fitness Prediction. Deep Mutational Scanning (DMS) datasets (icml2022Tranception, ; nguyen2024evo, ) introduce exhaustive mutations into protein coding sequences (CDS) and report fitness scores that reflect functional efficacy. For both bacterial and human proteins, we compute Spearman’s rank correlation (SRCC) between model scores and experimental measurements (Table 3). The protein-specialist ESM-2 sets the upper bound, yet Life-Code surpasses all DNA-based baselines including GenSLM (ijhpca2023genslms, ) and the 7-B-parameter EVO model—despite using fewer parameters. We attribute this advantage to Life-Code’s unified pre-training on DNA, RNA, and amino acids, which captures translation constraints and folding cues along the central-dogma pathway, enabling strong generalization from nucleotide space to protein fitness without any task-specific fine-tuning.
@@
Table 6: Comparison on Species classification. Top-1 accuracy (%) for 5-way classification with scaling up the sequence length, where OOM denotes out-of-memory on the single A100-80G.
@@
Figure 8: Model accuracy vs. throughput in K tokens/s for different hybrid layer ratios. A shift from 11:1 to 5:1 slightly slows throughput with minimal accuracy loss; 3:1 exceeds memory limits.
3.5 Comparison on Multi-omics Tasks
We conduct further comparisons with proposed biological foundation models (EVO (nguyen2024evo, ) and LucaOne (he2024lucaone, )) with multi-omics tasks. Following LucaOne, we first evaluate the non-coding RNA and protein interactions (ncRPI) task using SFT evaluation, which aims to predict the interaction strengths between ncRNAs and proteins. As shown in Table 5, Life-Code could take cDNA and Amino Acid sequences in the unified input format and achieve a competitive result as the supervised LucaOne, which should process two inputs with separate modules. Then, we evaluate the modeling of transcription and translation processes with binary classification tasks between cDNA and protein sequences, where Life-Code achieves the best accuracy over existing models.
3.6 Empirical Analysis
Figure 8 shows that a pure RNN stack (12:0) is fastest ( $\sim$45 K token/s) but sacrifices 2–3 pp of accuracy. Injecting just one MHSA layer (11:1) restores nearly all lost performance—DNA 95 to 97 %, Protein 32 to 34 %, CD 83 to 86 %—while reducing throughput by only $\sim$9 %. A 5:1 mix boosts accuracy further with a modest extra cost, whereas 3:1 exhausts memory. Thus, adding a small fraction of MHSA markedly improves downstream accuracy without a large computational penalty.
Table 7: Ablation Study of Life-Code tokenizer, encoder, and pre-training tasks with DNA, protein, and central dogma (CD) tasks. Note that the blue background denotes the selected setups.
@@ 4 Related Works
Recent DNA foundation models have evolved swiftly from DNABERT ji2021dnabert , which first treated genomic 6-mers as linguistic tokens, to DNABERT-2 zhou2023dnabert2 and the large-scale Nucleotide Transformer NM2023NucleotideTrans . To tame quadratic attention on longer sequences, HyenaDNA nguyen2024hyenadna replaces full self-attention with efficient hierarchical kernels. Other variants embed biological priors—Caduceus schiff2024caduceus for long-range structure and MxDNA nips2024MXDNA for adaptive tokenization—illustrating a broad shift toward specialized, scalable architectures for genomics.
@@
AlphaFold jumper2021alphafold ; jumper2021alphafold2 ignited the protein-LM wave, inspiring models like ESM-2 lin2022ESM2 . RNA efforts followed, e.g., RNA-FM shen2024rnafm for 3D structure prediction. While these biomolecule-specific architectures excel individually, recent work now fuses DNA, codon, and protein signals prakash2024bridging ; bioinfo2024genomic ; BioInfo2023NTA ; nmil2024CaLM , delivering notable downstream gains Ren2024COMET and pointing toward unified multi-omics models.
@@
Recent studies aim to bridge the gap between genomic and protein sequences, pushing beyond single-modality tasks. CD-GPT zhu2024CDGPT explicitly connects DNA, RNA, and proteins through the central dogma, while BSM xiang2024BSM highlights the potential for small but effective models covering genes and proteins simultaneously. EVO nguyen2024evo similarly integrates molecular and genome-scale data with a generalized sequence-modeling approach. Closely related is LucaOne he2024lucaone , which advocates a unified nucleic acid and protein language for biological representation.
@@ 5 Conclusion and Limitations
Figure A1: Data Collection Pipeline. We collect the reference sequences from the NCBI RefSeq database as the DNA dataset and collect the cDNA of coding sequences with the corresponding amino acids from the GenBank and UniRef50 databases to build up our DNA-AA pairing datasets.
Table A1: Configuration of pre-training datasets. As for the DNA dataset, we collect DNA/RNA sequences from the NCBI RefSeq database. As for the DNA-AA pairing dataset, we collect the cDNA (coding sequences) and its corresponding Amino Acids (or using translation) from the GenBank database and also collect Amino Acids and their corresponding cDNA sequences using reverse translation from the UniRef50 database.
@@
cDNA-AA pair 16,786,593 GenBank
@@
mRNA-AA pair 3.908,074 GenBank
@@ A.1 Pre-training Dataset
We collect two datasets for pre-training of Life-Code models, i.e., a pure DNA dataset and a DNA-AA pairing dataset, with the collection process shown in Figure A1. As for the DNA dataset, we collect the reference sequences (RefSeq) of multiple species to ensure generalization abilities from the database of the National Center for Biotechnology Information (NCBI) at https://www.ncbi.nlm.nih.gov, following the Multi-species Genomes [^1] Multi-species Genomes are originally provided in https://huggingface.co/datasets/InstaDeepAI/multi_species_genomes, which is further extended by DNABERT-2 in https://github.com/MAGICS-LAB/DNABERT_2 provided by Nucleotide Transformer [8] and DNABERT2 [64]. As for the DNA-AA pairing dataset, we collect the cDNA of coding sequences (CDS) and its corresponding Amino Acids (AA) in the GenBank database at https://www.ncbi.nlm.nih.gov/genbank, which aims to model the transcription and translation processes of the central dogma. We also collect some Amino Acids in the UniRef50 database following LucaOne [13] and obtain their corresponding cDNA by reverse translation with online tools. We provide detailed information for used datasets in Table A1.
@@ A.2 Life-Code Tokenizer
There are two vocabularies used in Life-Code. The unified vocabulary only uses 4 nucleotides {A, T/U, C, G} of nucleic acid with 5 special tokens, including “[U]”/“[UNK]”, “[PAD]”, “[CLS]”, “[SEP]”, and “[MASK]” for unknown nucleotides, padding tokens, the class token, separator tokens, and masking tokens. Meanwhile, the Life-Code can also use the codon vocabulary (i.e., the 3-mer of 4 nucleotides that constructs 64 codon tokens), which could be merged into 20 amino acids of protein (20 uppercase letters excluding “B”, “J”, “O”, “U”, “X”, and “Z”). It can only be applied when the length of an input sequence is multiples of 3, i.e., the cDNA of amino acids or matured mRNA (CDS). The pre-trained protein language model employs the amino acid vocabulary of ESM-2 [26].
Table A2: Configuration of the network designs and pre-training settings for Life-Code models. GDN blocks denote the Gated DeltaNet block with linear complexity, while Attention blocks denote the self-attention block with FLASH-Attention implementation.
@@
Attention heads 0 24
Parameters 8M 350/500M
@@ A.3 Life-Code Encoder
As shown in Table A2, the Life-Code encoder has 24 layers in total with the embedding dim of 1024 with the following designs: (1) Mixture of GDN blocks (DeltaNet [60]) and multi-head self-attention (MHSA) blocks as a hybrid model, especially every 11 GDN blocks followed by a self-attention block like MiniMax-01 [32], which could utilize the complementary properties of GDN and MHSA while maintaining efficiency. Notice that the DeltaNet variants [59] introduce short-term depth-wise convolution after the query/key/value projections in the token mixing block, which requires slightly more parameters than the standard Transformer block. (2) The model macro design employs pre-norm [55] with RMSNorm [61], Layer Scale [52], Rotary Position Embedding (RoPE) [51], SwiGLU [53], and FlashAttention implementations to facilitate training large-scale models stably with long sequences. (3) During pre-training, we apply the packing strategy [57] to build up a long sequence with several CDS, which compromises the gap between different lengths of the reference sequence and the coding sequences, as shown in Figure 3.
@@
As shown in Figure 5, we further pre-train the Life-Code tokenizer and Encoder with three tasks in Eq. 5 for 1M steps with a batch size of 256 and a basic learning rate of $1\times 10^{-4}$. We adopt 15% random masking in BERT for Masked DNA Reconstruction and the 3-mer span masking for CDS-to-Amino-Acid Translation. As for Knowledge Distillation from a pre-trained Protein LM, we adopted ESM2-650M (esm2_t33_650M_UR50D[26] and a protein decoder with the output dimension of 1280. During the warmup periods, the maximum sequence length is 1024, with a linear warmup of the learning rate for 10 iterations. After that, the maximum sequence length is set to 4k with the learning rate adjusted by the Cosine Annealing scheduler (decay to $1\times 10^{-6}$). We utilize 8 Nvidia A100-80G GPUs with a per-GPU batch size of 2 and a gradient accumulation time of 16.

### v3

@@ 3.2 Comparison on Genomic Benchmarks
As DNA sequences are the foundation of the central dogma, we primarily compare with the latest DNA foundation models with the SFT evaluation protocol proposed by (nips2024hyenadna; iclr2024dnabert2), following benchmark evaluations for fair comparison. We consider three types of DNA model architectures: (a) classical BERT encoders, including DNABERT (BioInfo2021dnabert), DNABERT2 (iclr2024dnabert2), GENA-LM (fishman2023GenaLM), Nucleotide Transformer (NT) (NM2023NucleotideTrans), with different tokenizers like VQDNA (icml2024vqdna) and MxDNA (nips2024MXDNA); (b) RNN-based linear attentions including HyenaDNA (nips2024hyenadna) and Caduceus (icml2024caduceus); (c) CNN model ConvNova (iclr2025convnova).
@@
We first compare Life-Code with 8 evaluation tasks in the Genomic Benchmark (BMC2023genomicbenchmark), which we adopted results from GenBench implementations (liu2024genbench), and all compared DNA models are fully SFT and reported the top-1 accuracy. Table 1 demonstrates that Life-Code surpasses previous models since its hybrid designs combine the advantages.
@@
GUE Benchmarks.
We then evaluate Life-Code with 24 long-range genomic tasks on the GUE benchmark (iclr2024dnabert2), as shown in Table 2, where 7 widely adopted practical genomic tasks are conducted, including Epigenetic Mark Prediction (EMP) for Yeast, Transcription Factor Prediction on Mouse and Human genome, Promoter Detection, Core Promoter Detection, Splice Site Prediction, and Covid Variants Classification. All results are adopted from DNABERT2 or each paper with their official SFT setups.
3.3 Comparison on RNA Splicing Site Prediction
Since precursor messenger RNAs (pre-mRNAs) splicing is a crucial process in eukaryotic gene expression, we then evaluate Life-Code on Splicing Site Prediction tasks as binary classification tasks. We consider two typical datasets with the SFT evaluation protocol. As for the SpliceAI dataset (JAGANATHAN2019SpliceAI), we evaluate DNA models with top-1 AUC scores and find that Life-Code consistently outperforms popular SpliceAI and DNA models, indicating its great ability to model the central dogma. As for Spliceator dataset (BMC2021spliceator), we then consider RNA splicing methods (Spliceator, DDSP (naito2019DDSP), RINALMo (Penic2024RiNALMo)) and RNA language model, RNA-FM (shen2024rnafm), with RNA data of multi-species. As shown in Table 4, our Life-Code achieves the best F1 scores.
@@ 3.3 Comparison on RNA Splicing Site Prediction
Table 4: Splicing Prediction with two splicing datasets from SpliceAI and Spliceator. The top-1 AUC score is reported.
@@ 3.4 Comparison on Protein Fitness Prediction
Following EVO (nguyen2024evo), we evaluate foundation models in a strict zero-shot setting on Protein Fitness Prediction. Deep Mutational Scanning (DMS) datasets (icml2022Tranception; nguyen2024evo) introduce exhaustive mutations into protein coding sequences (CDS) and report fitness scores that reflect functional efficacy. For both bacterial and human proteins, we compute Spearman’s rank correlation (SRCC) between model scores and experimental measurements (Table 3). The protein-specialist ESM-2 sets the upper bound, yet Life-Code surpasses all DNA-based baselines including GenSLM (ijhpca2023genslms) and the 7-B-parameter EVO model—despite using fewer parameters. We attribute this advantage to Life-Code’s unified pre-training on DNA, RNA, and amino acids, which captures translation constraints and folding cues along the central-dogma pathway, enabling strong generalization from nucleotide space to protein fitness without fine-tuning.
3.5 Comparison on Multi-omics Tasks
We conduct further comparisons with proposed biological foundation models (EVO (nguyen2024evo) and LucaOne (he2024lucaone)) with multi-omics tasks. Following LucaOne, we first evaluate the non-coding RNA and protein interactions (ncRPI) task using SFT evaluation, which aims to predict the interaction strengths between non-coding RNAs (ncRNAs) and proteins. As shown in Table 5, Life-Code can accept cDNA and Amino Acid sequences in a unified input format and achieve a competitive result, comparable to the supervised LucaOne, which requires separate modules for each input. Then, we evaluate the modeling of transcription and translation processes using binary classification tasks between cDNA and protein sequences, where Life-Code achieves the best accuracy compared to existing models. We can see that the modalities enhance each other.
@@ 3.5 Comparison on Multi-omics Tasks
Table 6: Comparison on Species classification. Top-1 accuracy (%) for 5-way classification with scaling up the sequence length, where OOM denotes out-of-memory.
Figure 8: Model accuracy vs. throughput in K tokens/s for different hybrid layer ratios. A shift from 11:1 to 5:1 slightly slows throughput with minimal accuracy loss.
3.6 Empirical Analysis of Hybrid Design
Figure 8 shows that a pure RNN stack (12:0) is fastest ( $\sim$45 K token/s) but sacrifices 2–3 pp of accuracy. Injecting just one MHSA layer (11:1) restores nearly all lost performance—DNA 95 to 97 %, Protein 32 to 34 %, CD 83 to 86 %—while reducing throughput by only $\sim$9 %. A 5:1 mix boosts accuracy further with a modest extra cost, whereas 3:1 exhausts memory. Thus, adding a small fraction of MHSA markedly improves downstream efficiency.
@@ 3.6 Empirical Analysis of Hybrid Design
Table 7: Ablation Study of Life-Code and pre-training tasks with DNA, protein, and central dogma (CD) tasks. Note that the blue background denotes the selected setups.
@@ 4 Related Works
Recent DNA foundation models have evolved swiftly from DNABERT (ji2021dnabert), which first treated genomic 6-mers as linguistic tokens, to DNABERT-2 (zhou2023dnabert2) and the large-scale Nucleotide Transformer (NM2023NucleotideTrans). To tame quadratic attention on longer sequences, HyenaDNA (nguyen2024hyenadna) replaces full self-attention with efficient hierarchical kernels. Other variants embed biological priors—Caduceus (schiff2024caduceus) for long-range structure and MxDNA (nips2024MXDNA) for adaptive tokenizer.
@@
AlphaFold (jumper2021alphafold; jumper2021alphafold2) ignited the protein-LM wave, inspiring models like ESM-2 (lin2022ESM2). RNA efforts followed, e.g., RNA-FM (shen2024rnafm) for 3D structure prediction. While these biomolecule-specific architectures excel individually, recent work now fuses DNA, codon, and protein signals (prakash2024bridging; bioinfo2024genomic; BioInfo2023NTA; nmil2024CaLM), delivering notable downstream gains (Ren2024COMET) and toward unified models.
@@
Recent studies aim to bridge the gap between genomic and protein sequences, pushing beyond single-modality tasks. CD-GPT (zhu2024CDGPT) explicitly connects DNA, RNA, and proteins through the central dogma, while BSM (xiang2024BSM) highlights the potential for small but effective models covering genes and proteins simultaneously. EVO (nguyen2024evo) similarly integrates molecular and genome-scale data with a generalized sequence-modeling approach. Closely related is LucaOne (he2024lucaone), which advocates a unified nucleic acid and protein language for biological representation.
@@ 5 Conclusion and Limitations
Figure A1: Data Collection Pipeline. We collect the reference sequences from the NCBI RefSeq database as the DNA dataset and collect the cDNA of coding sequences with the corresponding amino acids from the GenBank and UniRef50 databases to build our DNA-AA pairing datasets.
@@
cDNA-AA pair 49,453,295 GenBank
@@
mRNA-AA pair 25,468,310 GenBank
Table A1: Configuration of pre-training datasets. As for the DNA dataset, we collect DNA/RNA sequences from the NCBI RefSeq database. As for the DNA-AA pairing dataset, we collect the cDNA (coding sequences) and their corresponding Amino Acids (or using translation) from the GenBank database, and also collect Amino Acids and their corresponding cDNA sequences using reverse translation from the UniRef50 database.
@@ A.1 Pre-training Dataset
We collect two datasets for pre-training Life-Code models, i.e., a pure DNA dataset and a DNA-AA pairing dataset, as shown in Figure A1. As for the DNA dataset, we collect the reference sequences (RefSeq) of multiple species to ensure generalization abilities from the database of the National Center for Biotechnology Information (NCBI) at https://www.ncbi.nlm.nih.gov, following the Multi-species Genomes [^1] Multi-species Genomes are originally provided in https://huggingface.co/datasets/InstaDeepAI/multi˙species˙genomes, which is further extended by DNABERT-2 in https://github.com/MAGICS-LAB/DNABERT˙2 provided by Nucleotide Transformer (NM2023NucleotideTrans) and DNABERT2 (iclr2024dnabert2). It includes around 135 species in 7 categories with 35.4B nucleotide bases in total. As for the DNA-AA pairing dataset, we collect the cDNA of coding sequences (CDS) and their corresponding Amino Acids (AA) in the GenBank database at https://www.ncbi.nlm.nih.gov/genbank, which aims to model the transcription and translation processes of the central dogma. We also collect some Amino Acids in the UniRef50 database following LucaOne (he2024lucaone) and obtain their corresponding cDNA by reverse translation with online tools to build cDNA-AA pairs. It contains 92M sequences in total. We provide detailed information for the used datasets in Table A1.
@@ A.2 Life-Code Tokenizer
There are two vocabularies used in Life-Code. The unified vocabulary only uses 4 nucleotides {A, T/U, C, G} of nucleic acid with 5 special tokens, including “[U]”/“[UNK]”, “[PAD]”, “[CLS]”, “[SEP]”, and “[MASK]” for unknown nucleotides, padding tokens, the class token, separator tokens, and masking tokens. Meanwhile, the Life-Code can also use the codon vocabulary (i.e., the 3-mer of 4 nucleotides that constructs 64 codon tokens), which could be merged into 20 amino acids of protein (20 uppercase letters excluding “B”, “J”, “O”, “U”, “X”, and “Z”). It can only be applied when the length of an input sequence is a multiple of 3, i.e., the cDNA of amino acids or matured mRNA (CDS). The pre-trained protein language model employs the amino acid vocabulary of ESM-2 (lin2022ESM2).
@@
Attention heads 0 16
Parameters 8M 305/500M
@@
Table A2: Configuration of the network designs and pre-training settings for Life-Code models. GDN blocks denote the Gated DeltaNet (yang2024gateddelta) block with linear complexity, while Attention blocks denote the self-attention block with FLASH-Attention implementation (Dao2022FlashAttention).
@@ A.3 Life-Code Encoder
As shown in Table A2, the Life-Code encoder has 24 layers in total with the embedding dim of 1024 with the following designs: (1) Mixture of GDN blocks (DeltaNet (yang2025deltanet)) and multi-head self-attention (MHSA) blocks with the head dim of 64 as a hybrid model, especially every 11 GDN blocks followed by a self-attention block like MiniMax-01 (MiniMax2025MiniMax01), which could utilize the complementary properties of GDN and MHSA while maintaining efficiency. Notice that the DeltaNet variants (yang2024gateddelta) introduce short-term depth-wise convolution after the query/key/value projections in the token mixing block, which requires slightly more parameters than the standard Transformer block. (2) The model macro design employs pre-norm (acl2019PreNorm) with RMSNorm (Zhang2019RMSNorm), Layer Scale (iccv2021CaiT), Rotary Position Embedding (RoPE) (Su2021RoFormer), SwiGLU (Touvron2023LLaMA), and FlashAttention implementations to facilitate training large-scale models stably with long sequences. (3) During pre-training, we apply the packing strategy (Warner2024ModernBERT) to build up a long sequence with several CDS, which compromises the gap between different lengths of the reference sequence and the coding sequences, as shown in Figure 3.
@@
As shown in Figure 7, we further pre-train the Life-Code tokenizer and Encoder with three tasks in Eq. 5 for 1M steps with a batch size of 256 and a basic learning rate of $1\times 10^{-4}$. We adopt 15% random masking in BERT for Masked DNA Reconstruction and the 3-mer span masking for CDS-to-Amino-Acid Translation. As for Knowledge Distillation from a pre-trained Protein LM, we adopted ESM2-650M (esm2_t33_650M_UR50D(lin2022ESM2) and a protein decoder with the output dimension of 1280. During the warmup periods, the maximum sequence length is 1024, with a linear warmup of the learning rate for 10 iterations. After that, the maximum sequence length is set to 4k with the learning rate adjusted by the Cosine Annealing scheduler (decay to $1\times 10^{-6}$). We utilize 8 Nvidia A100-80G GPUs with a per-GPU batch size of 2 and a gradient accumulation time of 16.